5 -letter "words'' are formed using the letters A, B, C, D, E, F, G.
(Note that these "words" do not have to be real words and letters can be reused unless told otherwise)
How many such words are possible for each of the following conditions?
(a) No condition is imposed.
Your answer is :
(b) No letter can be repeated in a word.
Your answer is :
(c) Each word must begin with the letter A.
Your answer is :
(d) The letter C must be at the end.
Your answer is :
(e) The second letter must be a vowel.
Your answer is :

Answer by Boreal(15235)

a. 7 ways to choose the first*7 for the second...*7 for the fifth.
That is 16807 ways.
b. that is 7*6*5*4*3=2520 ways
c. that is 1*7*7*7*7=2401 ways
d. That is 7*7*7*7*1=2401 ways
e. that is 7*2*7*7*7=4802 ways